2 AlBr3 + 3 Cl2 → 2 AlCl3 + 3 Br2

If a reaction occurs using 3.75 moles AlBr3, how many moles of Brwould be produced? 


1
Expert's answer
2022-05-16T22:01:01-0400

2AlBr3 + 3Cl2 = 2AlCl3 + 3Br2

According to the above reaction, 3mol Br2 is formed when 2mol AlBr3 is used.

Now we find how many moles of Br2 are formed from 3.75 moles of AlBr3.

2mol AlBr3 —— 3 mol Br2

3.75mol AlBr3 —— x mol Br2

x = 3.75 * 3 / 2 = 5.625 mol

Answer: 5.625 mol Br2